OpenStudy (anonymous):

Write an equation of the line with a y-intercept of -4 that is parallel to 6x-2y=13

OpenStudy (anonymous):

convert to slope-intercept form Then write the same slope and the y-intercept. Find b. Done.

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

Do you know what slope-intercept form looks like?

OpenStudy (anonymous):

No I don't, can you please show me

OpenStudy (anonymous):

Y slope is y=x+b

OpenStudy (anonymous):

so all you do is subtract and then divide to get y by it self

OpenStudy (anonymous):

Actually that is a wrong representation. Misleading and unclear information. Let me take it from here.

OpenStudy (anonymous):

sorry its y=mx+b

OpenStudy (anonymous):

Slope intercept form is y = mx + b. Your coordinate is (x,y) Substitute it in the equation where x and y is located. b is the vertical translation I believe. Clarify. And m is your slope.

OpenStudy (anonymous):

Yes Music Your equation is correct. But she needs to learn.

OpenStudy (anonymous):

I know what my slope is, but I need a parallel line. Every time I try to write one up, it isn't parallel.

OpenStudy (anonymous):

@dumbcow @UnkleRhaukus @.Sam.

OpenStudy (anonymous):

What is your slope?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Nope why dont you do the steps and ill show you what i did wrong.

OpenStudy (anonymous):

What are you talking about?

OpenStudy (anonymous):

6 is not the slope.

OpenStudy (anonymous):

I hope you are writing the steps music Better show her how its done

OpenStudy (anonymous):

First of all KimberlyAlice, all you do to get a parallel line is to use the same slope, because lines that are parallel have the same slope. Now to get the slope of the equation it gives you you have to change it into a slope-intercept equation, which is y=mx+b. So change 6x-2y=13 into y=mx.b

OpenStudy (anonymous):

No your teachers have perhaps been teaching you right. You are misinterpreting them. Your equation is in Ax + By = C you need y = mx + b

OpenStudy (anonymous):

you need to find m by isolating y

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

do you have it now then?